# Optimal discrimination of quantum sequences

### v1

@@ Appendix C Proof of Lemma 5
For each $i\in[k]$, let $\{u_{1}^{i},\ldots,u_{s_{i}}^{i}\}$ be a basis of $\mathcal{W}_{i}$, and extend it to a basis $\{u_{1}^{i},\ldots,u_{s_{i}}^{i},v_{1}^{i},\ldots,v_{t_{i}}^{i}\}$ of $\mathcal{V}_{i}$. First, express each $a_{i}$ in terms of the basis of $\mathcal{V}_{i}$:
@@
where $b_{i}\in\mathcal{W}_{i}$. Since $a_{i}\notin\mathcal{W}_{i}$, there exists at least one $j_{i}$ such that $\lambda_{j_{i}}^{i}\neq 0$. Consider the tensor product $a_{1}\otimes\cdots\otimes a_{k}$. Its expansion includes the term:
@@
This term is non-zero because each $\lambda_{j_{i}}^{i}\neq 0$. Observe that $v_{j_{1}}^{1}\otimes\cdots\otimes v_{j_{k}}^{k}\notin\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$, as each $v_{j_{i}}^{i}\notin\mathcal{W}_{i}$. Finally, note that $a_{1}\otimes\cdots\otimes a_{k}\notin\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$, because its expansion contains a non-zero term that is not in $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$. ∎
@@
Let $\mathcal{V}_{i}$ be vector spaces and $B_{i}=\{b_{1}^{i},\ldots,b_{\ell_{i}}^{i}\}\subseteq\mathcal{V}_{i}$ for $i\in[k]$. Let $\mathcal{W}_{i}=\text{span}(B_{i})$ be the subspaces spanned by $B_{i}$. Then
@@
The eigenvectors corresponding to non-zero eigenvalues of $\rho_{y_{1}}^{1}\otimes\cdots\otimes\rho_{y_{k}}^{k}$ are
@@
Appendix D Proof of Theorem 9
@@
We expand both sides of Equation (33) in terms of the kernels. The right side can be written as
@@
The left side of Equation (33) can be expanded as follows
@@
Now for each $k$-tuple $(y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})$, consider $\operatorname{ker}(\rho^{1}_{y_{1}})\otimes\mathbb{C}^{d_{2}}\otimes\cdots\otimes\mathbb{C}^{d_{k}}+\cdots+\mathbb{C}^{d_{1}}\otimes\cdots\otimes\mathbb{C}^{d_{k-1}}\otimes\operatorname{ker}(\rho^{k}_{y_{k}})$. This sum contains $\operatorname{ker}(\rho^{1}_{y_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{y_{k}})$ and thus contains $\bigcap\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right)$ where $j_{i}\neq x_{i}$ for all $i\in[k]$. Therefore, we conclude:
@@
For the converse, we will show that the following relation holds
@@
where for a subspace $\mathcal{V}$, $\mathcal{V}^{\perp}$ denotes its orthogonal complement. First, we express the right side of the above relation as follows:
@@
where $\operatorname{Im}(\rho)$ denotes the image of an operator $\rho$. In the second equality, we have used the fact that the orthogonal complement of the intersection of subspaces is equal to the sum of the orthogonal complement of individual subspaces: $(\mathcal{V}_{1}\cap\cdots\cap\mathcal{V}_{n})^{\perp}=\mathcal{V}_{1}^{\perp}+\cdots+\mathcal{V}_{n}^{\perp}$ where $\mathcal{V}_{i}$’s are subspaces. This identity also allows us to write the left side of Equation (39) as a sum of subspaces
@@
Each term in the above sum is of the form $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{n}$ where $\mathcal{W}_{i}$ is either $\mathcal{V}_{i}$ or $\mathcal{V}_{i}^{\perp}$, and there is at least one index $a\in[n]$ for which $\mathcal{W}_{a}=\mathcal{V}_{a}^{\perp}$. Therefore, each term on the right side of Equation (41) is a sum of terms like $(\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}})$. That is,
@@
Therefore, the pure tensor, which was assumed to be an element of $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$ is shown to belong to
@@
Now observe that $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$ is spanned by pure tensors and since
@@
is sum of spaces of the form $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$, it is spanned by pure tensors as well. The span of these pure tensors form a subspace in
@@
Appendix E Semidefinite program of unambiguous discrimination of sequences
The SDPs of the ensembles and the sequences follow straightforwardly from the SDP given in Equation (8) in the main text. However, for the reader’s convenience, we write them here explicitly.
The SDP for the optimal probability of unambiguous discrimination of the states of ensemble $\mathcal{E}^{i}$ is the following
Primal problem maximize: $\sum_{j=1}^{\ell_{i}}\eta_{j}^{i}\operatorname{Tr}(\rho_{j}^{i}\Theta_{j}^{i}\Delta_{j}^{i}{\Theta_{j}^{i}}^{\dagger})$ subject to: $\sum_{j=1}^{\ell_{i}}\Theta_{j}^{i}\Delta_{j}^{i}\Theta_{j}^{i\dagger}\preceq\mathbb{1},$ $\Delta_{j}^{i}\succeq 0,\quad\forall j\in[\ell_{i}]$              Dual problem minimize: $\operatorname{Tr}(Z_{i})$ subject to: ${\Theta_{j}^{i}}^{\dagger}\left(Z_{i}-\eta_{j}^{i}\rho_{j}^{i}\right)\Theta_{j}^{i}\succeq 0\quad\forall j\in[\ell_{i}],$ $Z_{i}\succeq 0.$
Here $\Theta_{j}^{i}$ is an $d_{i}\times r_{j}^{i}$ matrix whose columns form an arbitrary orthonormal basis for $S_{j}^{i}$ (of dimension $r_{j}^{i}$).
@@ Appendix E Semidefinite program of unambiguous discrimination of sequences
The SDP for its optimal unambiguous discrimination is the following
@@
Here the matrices $\Theta(x_{1},...,x_{k})$ and $\Delta(x_{1},...,x_{k})$ are $\ell$ in number, one for each $k$-tuple $(x_{1},...,x_{k})$.

### v2

@@ IV.1 Conditions for unambiguous discrimination of quantum sequences
For each $i\in[k]$, let $\{u_{1}^{i},\ldots,u_{s_{i}}^{i}\}$ be a basis of $\mathcal{W}_{i}$, and extend it to a basis $\{u_{1}^{i},\ldots,u_{s_{i}}^{i},v_{1}^{i},\ldots,v_{t_{i}}^{i}\}$ of $\mathcal{V}_{i}$. First, express each $a_{i}$ in terms of the basis of $\mathcal{V}_{i}$,
@@
where $b_{i}\in\mathcal{W}_{i}$. Since $a_{i}\notin\mathcal{W}_{i}$, there exists at least one $j_{i}$ such that $\lambda_{j_{i}}^{i}\neq 0$. Consider the tensor product $a_{1}\otimes\cdots\otimes a_{k}$. Its expansion includes the term
@@
This term is nonzero because each $\lambda_{j_{i}}^{i}\neq 0$. Observe that $v_{j_{1}}^{1}\otimes\cdots\otimes v_{j_{k}}^{k}\notin\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$, as each $v_{j_{i}}^{i}\notin\mathcal{W}_{i}$. Finally, note that $a_{1}\otimes\cdots\otimes a_{k}\notin\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$ because its expansion contains a nonzero term that is not in $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$. ∎
@@
Let $\mathcal{V}_{i}$ be vector spaces and $B_{i}=\{b_{1}^{i},\ldots,b_{\ell_{i}}^{i}\}\subseteq\mathcal{V}_{i}$ for $i\in[k]$. Let $\mathcal{W}_{i}=\text{span}(B_{i})$ be the subspaces spanned by $B_{i}$. Then,
@@
The eigenvectors corresponding to nonzero eigenvalues of $\rho_{y_{1}}^{1}\otimes\cdots\otimes\rho_{y_{k}}^{k}$ are
@@
The following theorem is immediate from Lemmas 4 and 5.
Theorem 8 (Condition for unambiguous discrimination of quantum sequences).
IV.2 Proof of Theorem 1
We now present the proof of Theorem 1 for the unambiguous case. The unambiguous discrimination problem of mixed quantum states can be cast as a semidefinite program (SDP) [35]. Let us first present this SDP for an ensemble of states $\mathcal{E}=(Q,q)$, where $Q=\{\sigma_{i}\}_{i=1}^{N}$ is a set of density operators acting on $\mathbb{C}^{d}$ and $q=(q_{1},\ldots,q_{N})$ is the vector of a priori probabilities of states of $Q$. Without loss of generality, we can assume that the eigenvectors of the states of $Q$ that correspond to nonzero eigenvalues span $\mathbb{C}^{d}$. This is analogous to assuming that for unambiguous discrimination of pure states, we can take the dimension of the Hilbert space they live in to equal the dimension of the space spanned by them. We define $\widetilde{Q_{i}}$ as the intersection of all kernels $\mathcal{K}_{j}$ of $\sigma_{j}$, excluding $\mathcal{K}_{i}$. This can be expressed mathematically as $\widetilde{Q_{i}}=\cap_{j=1,j\neq i}^{N}\mathcal{K}_{j}$. We also introduce $\Theta_{i}$, a $d\times r_{i}$ matrix whose columns form an orthonormal basis for $\widetilde{Q_{i}}$, where $r_{i}$ is the dimension of $\widetilde{Q_{i}}$. With these definitions in place, we can formulate the SDP for determining the optimal probability of unambiguous discrimination of the ensemble $\mathcal{E}$ as follows:
Primal problem maximize: $\sum_{i=1}^{N}q_{i}\operatorname{Tr}(\sigma_{i}\Theta_{i}\Delta_{i}\Theta_{i}^{\dagger})$ subject to: $\sum_{i=1}^{N}\Theta_{i}\Delta_{i}\Theta_{i}^{\dagger}\preceq\mathbb{1},$ $\Delta_{i}\succeq 0,\quad\forall i\in[N]$              Dual problem minimize: $\operatorname{Tr}(Z)$ subject to: $\Theta_{i}^{\dagger}\left(Z-q_{i}\sigma_{i}\right)\Theta_{i}\succeq 0\quad\forall i\in[N],$ $Z\succeq 0$ (28)
where $\Delta_{i}$ is an $r_{i}\times r_{i}$ matrix for each $i\in[N]$. The optimization variables in this formulation are the $N$ matrices denoted by $\Delta_{i}$. Our proof strategy will be to use Slater’s theorem, which states that if the primal problem is convex and strict feasibility holds, then the duality gap is zero and the primal and dual optimal values are equal [36]. We will first assume primal and dual optimal variables for the ensemble SDPs. Then we will use these variables to construct primal and dual feasible solutions for the sequence SDP. Finally, we will show that these variables make the primal and dual value equal. Applying Slater’s theorem to the sequence SDP, we will conclude that these must be the optimal solutions.
Proof.
Consider the ensemble
where $\rho_{j}^{i}$ are $d_{i}\times d_{i}$ density operators and $j\in[\ell_{i}]$. Let $S^{i}_{j}=S(\mathcal{E}^{i})\setminus\{\rho^{i}_{j}\}$ for some $j\in[\ell_{i}]$ and $\widetilde{S^{i}_{j}}$ be the intersection of the kernels of all the density matrices of $\mathcal{E}^{i}$, except for $\rho^{i}_{j}$. That is, $\widetilde{S^{i}_{j}}=\cap_{t=1,t\neq j}^{\ell_{i}}\mathcal{K}_{t}^{i}$, where $\mathcal{K}_{t}^{i}$ is the kernel of $\rho_{t}^{i}$. Let $\Theta_{j}^{i}$ be a $d_{i}\times r_{j}^{i}$ matrix whose columns form an arbitrary orthonormal basis for $\widetilde{S^{i}_{j}}$ (which is of dimension $r_{j}^{i}$). The optimal probability of unambiguous discrimination of the elements of $\mathcal{E}^{i}$ is denoted by $p(\mathcal{E}^{i})$ and let $\Delta_{j}^{i}$ be the $r_{j}^{i}\times r_{j}^{i}$ matrices that achieve this optimum. In addition to being positive, these matrices satisfy $\sum_{j=1}^{\ell_{i}}\Theta_{j}^{i}\Delta_{j}^{i}{\Theta_{j}^{i}}^{\dagger}\preceq\mathbb{1}$. (See the Appendix for the SDPs of the individual ensembles and the sequence.) We denote the optimal dual variable for ensemble $\mathcal{E}^{i}$ by $Z_{i}$, which is also a positive matrix and satisfies ${\Theta_{j}^{i}}^{\dagger}\left(Z_{i}-\eta_{j}^{i}\rho_{j}^{i}\right)\Theta_{j}^{i}\succeq 0$. The optimal primal and dual variables satisfy
Now, we turn to the sequence ensemble given by Eq. (2) and consider the SDP of its unambiguous discrimination. Note that there are $\ell$ number of sequence states in this ensemble, and a state is labeled by the index $(x_{1},\ldots,x_{k})$, which is a $k$-tuple of indices where $x_{i}\in[\ell_{i}]$ for all $i\in[k]$. Let us denote by $\widetilde{S}(x_{1},\ldots,x_{k})$ the intersection of the kernels of all the states of $\mathcal{E}_{k}$ except for $\rho^{1}_{x_{1}}\otimes\cdots\otimes\rho^{k}_{x_{k}}$. That is,
where $\mathcal{K}(y_{1},\ldots,y_{k})$ is the kernel of $\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}}$. This subspace has a decomposition (see Theorem 9 in the Appendix) as follows
We first give a prescription to construct matrices $\Theta(x_{1},...,x_{k})$ whose columns form an orthonormal basis of $\widetilde{S}(x_{1},\ldots,x_{k})$. From Eq. (IV.2), it can be seen that these matrices are of size $\prod_{i=1}^{k}d_{i}\times\prod_{i=1}^{k}r^{i}_{x_{i}}(x_{i}\in[\ell_{i}])$ and a simple procedure to construct them is to take $\Theta(x_{1},...,x_{k})=\Theta_{x_{1}}^{1}\otimes\cdots\otimes\Theta_{x_{k}}^{k}$. For a feasible primal variable $\Delta(x_{1},...,x_{k})$, we take the tensor product of the optimal primal variables, $\Delta(x_{1},...,x_{k})=\Delta_{x_{1}}^{1}\otimes\cdots\otimes\Delta_{x_{k}}^{k}$. This operator is positive since it is the tensor product of positive operators. Also, the defined operators satisfy (using Lemma 3)
$\preceq$  $\;\mathbb{1}.$
Now consider the dual variable $Z=Z_{1}\otimes\cdots\otimes Z_{k}$, which is positive due to the positivity of the operators $Z_{i}$ for all $i\in[k]$. We also know that these variables satisfy ${\Theta_{x_{i}}^{i}}^{\dagger}Z_{i}\Theta_{x_{i}}^{i}\succeq{\Theta_{x_{i}}^{i}}^{\dagger}(\eta_{x_{i}}^{i}\rho_{x_{i}}^{i})\Theta_{x_{i}}^{i}$ for all $x_{i}\in[\ell_{i}]$ and $i\in[k]$. By taking tensor products for the indices $(x_{1},...,x_{k})$ we get, (using Lemma 3)
V Conclusions
Sequence discrimination problems arise in various quantum information processing tasks, where the objective is to determine the state of a sequence of finite length. The main question here is whether an optimal measurement is local or collective, i.e., whether it suffices to measure the individual members of the sequence or not. In this paper, we showed that as long as the members of the sequences are independently drawn from the same ensemble or even from different ensembles, the optimal measurement in both minimum-error and unambiguous discrimination paradigms is local and comprises fixed measurements on the individual components. It follows that if we give up the condition of independent selection of the members of a sequence, the optimal measurement could be either local and fixed, local and adaptive [18], or collective (partially or wholly) [19].
Our result also shows that independent sequences do not exhibit nonlocality in the sense some other unentangled systems do, as collective measurement is not necessary for optimal extraction of information from such sequences. However, in the context of quantum state exclusion, one may consider them nonlocal, for an entangled measurement is necessary, as demonstrated by the Pusey–Barrett–Rudolph result [37] and its generalizations [38].
A significant part of our proof relied on solving the SDP formulation of the problem using the strong duality theorem [36]. While SDP has been successfully used to solve various problems in quantum information theory [39, 40], we believe our approach could help to solve problems in many-body unentangled systems.
Our result could be immediately applied to settings of quantum key distribution protocols [41, 42] or similar ones. In these protocols, Alice sends a sequence of quantum states, selected independently from a known ensemble, to Bob, who measures each state individually. However, if Bob can store the incoming states in quantum memory, he may think about performing a collective measurement on the entire sequence to extract more information. Our result rules out this possibility, as these sequences are i.i.d.
A natural direction for future research involves doing away with the independence assumption. Can there be sequences of dependent quantum states whose optimal discrimination is still achievable by fixed, local measurements? A systematic study involving the set of states, the nature of the joint probability distribution for the states of a sequence, and the nature of optimal measurement to discriminate them will be interesting to investigate.
Acknowledgements.
Appendix A Proof of Theorem 9
@@
We expand both sides of Eq. (36) in terms of the kernels. The right side can be written as
@@
The left side of Eq. (36) can be expanded as follows:
@@
Now for each $k$-tuple $(y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})$, consider $\operatorname{ker}(\rho^{1}_{y_{1}})\otimes\mathbb{C}^{d_{2}}\otimes\cdots\otimes\mathbb{C}^{d_{k}}+\cdots+\mathbb{C}^{d_{1}}\otimes\cdots\otimes\mathbb{C}^{d_{k-1}}\otimes\operatorname{ker}(\rho^{k}_{y_{k}})$. This sum contains $\operatorname{ker}(\rho^{1}_{y_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{y_{k}})$ and thus contains $\bigcap\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right)$ where $j_{i}\neq x_{i}$ for all $i\in[k]$. Therefore, we conclude that
@@
For the converse, we will show that the following relation holds:
@@
where, for a subspace $\mathcal{V}$, $\mathcal{V}^{\perp}$ denotes its orthogonal complement. First, we express the right side of the above relation as follows:
@@
where $\operatorname{Im}(\rho)$ denotes the image of an operator $\rho$. In the second equality, we have used the fact that the orthogonal complement of the intersection of subspaces is equal to the sum of the orthogonal complement of individual subspaces: $(\mathcal{V}_{1}\cap\cdots\cap\mathcal{V}_{n})^{\perp}=\mathcal{V}_{1}^{\perp}+\cdots+\mathcal{V}_{n}^{\perp}$, where $\mathcal{V}_{i}$’s are subspaces. This identity also allows us to write the left side of Eq. (42) as a sum of subspaces,
@@
Each term in the above sum is of the form $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{n}$, where $\mathcal{W}_{i}$ is either $\mathcal{V}_{i}$ or $\mathcal{V}_{i}^{\perp}$, and there is at least one index $a\in[n]$ for which $\mathcal{W}_{a}=\mathcal{V}_{a}^{\perp}$. Therefore, each term on the right side of Eq. (44) is a sum of terms such as $(\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}})$. That is,
@@
Therefore, the pure tensor, which was assumed to be an element of $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$, is shown to belong to
@@
Now observe that $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$ is spanned by pure tensors and, since
@@
is the sum of spaces of the form $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$, it is spanned by pure tensors as well. The span of these pure tensors form a subspace in
@@
Appendix B Semidefinite program of unambiguous discrimination of sequences
The SDPs of the ensembles and the sequences follow straightforwardly from the SDP given in Eq. (28) in the main text. However, for the reader’s convenience, we write them here explicitly.
The SDP for the optimal probability of unambiguous discrimination of the states of ensemble $\mathcal{E}^{i}$ is the following:
Primal problem maximize: $\sum_{j=1}^{\ell_{i}}\eta_{j}^{i}\operatorname{Tr}(\rho_{j}^{i}\Theta_{j}^{i}\Delta_{j}^{i}{\Theta_{j}^{i}}^{\dagger})$ subject to: $\sum_{j=1}^{\ell_{i}}\Theta_{j}^{i}\Delta_{j}^{i}\Theta_{j}^{i\dagger}\preceq\mathbb{1},$ $\Delta_{j}^{i}\succeq 0,\quad\forall j\in[\ell_{i}]$              Dual problem minimize: $\operatorname{Tr}(Z_{i})$ subject to: ${\Theta_{j}^{i}}^{\dagger}\left(Z_{i}-\eta_{j}^{i}\rho_{j}^{i}\right)\Theta_{j}^{i}\succeq 0\quad\forall j\in[\ell_{i}],$ $Z_{i}\succeq 0.$ (54)
Here, $\Theta_{j}^{i}$ is an $d_{i}\times r_{j}^{i}$ matrix whose columns form an arbitrary orthonormal basis for $S_{j}^{i}$ (of dimension $r_{j}^{i}$).
@@ Appendix B Semidefinite program of unambiguous discrimination of sequences
The SDP for its optimal unambiguous discrimination is the following:
@@
Here, the matrices $\Theta(x_{1},...,x_{k})$ and $\Delta(x_{1},...,x_{k})$ are $\ell$ in number, with one for each $k$-tuple $(x_{1},...,x_{k})$.